# tttLRM: Test-Time Training for Long Context and Autoregressive 3D Reconstruction

### v1

@@ 2 Related Work
Multi-view 3D Reconstruction 3D reconstruction from images has been extensively studied in computer vision. Traditional methods such as structure-from-motion [41] or multi-view stereo (MVS) [15] focus on recovering 3D geometry. Deep learning has enabled feed-forward 3D reconstruction [61, 62, 25], which builds cost volumes using plane sweep for per-view depth estimation. Recently, learning-based MVS approaches [55, 59, 27, 53, 8, 26] directly estimate point clouds from input images and have been applied to camera pose estimation. Test3R [67] optimizes the network at test time in a self-supervised manner to improve 3D reconstruction. Concurrent work TTT3R [8] defines a gradient to update states for point cloud reconstruction. However, none of these methods can produce photo-realistic novel view synthesis.
@@ 4.4 Results
Autoregressive Reconstruction We demonstrate the autoregressive reconstruction capability of our model in the second row of LABEL:fig:teaser. With only 4 initial input views, the model already produces reasonable 3D Gaussian reconstructions; as additional views arrive (8 and 32 views), both the rendering quality and scene coverage progressively improve. Additional examples of autoregressive reconstruction are provided in the supplemental materials.
@@ 4.5 Ablation Study
Optimizer and Losses We use Muon optimizer for its stabilty and robustness. Table 5 shows that use Muon as opitmizer can bring better results even on low resolution setting. It will bring better results on longer sequence (e.g. high resolution and more input views). Also, incorporating depth and opacity as regularization can help reduce opaque Gaussians.
@@ B.1 Scene-level Training
We adopt a curriculum training strategy that progresses from low to high resolution, motivated by two main reasons. First, fast low-res pretraining enables the model to train with a large batch size with faster iteration time. Second, we found that even with pretrained TTT-LVSM [71] checkpoints at high-resolution (i.e. $960\times 640$), the model cannot predict reasonable Gaussians at the beginning iterations, leading to excessive GPU memory usage due to the rendering of a large number of Gaussians.
@@ Appendix C More results and Comparison
In Table 7, we provide a more detailed comparison including our model’s performance in autoregressive reconstruction mode, which also constantly outperforms Long-LRM and remains competitive with, or superior to, optimization-based baselines.
We also show results where we combine our method with a few additional optimization steps. It demonstrates that the reconstructed model can be further improved with minimal optimization cost, surpassing both purely optimization-based methods and the previous state-of-the-art feed-forward method, Long-LRM, under the same post-optimization setup.

### v2

@@ 2 Related Work
Multi-view 3D Reconstruction 3D reconstruction from images has been extensively studied in computer vision. Traditional methods such as structure-from-motion [40] or multi-view stereo (MVS) [14] focus on recovering 3D geometry. Deep learning has enabled feed-forward 3D reconstruction [60, 61], which builds cost volumes using plane sweep for per-view depth estimation. Recently, learning-based MVS approaches [54, 58, 26, 52, 7, 24] directly estimate point clouds from input images and have been applied to camera pose estimation. Test3R [66] optimizes the network at test time in a self-supervised manner to improve 3D reconstruction. Concurrent work TTT3R [7] defines a gradient to update states for point cloud reconstruction. However, none of these methods can produce photo-realistic novel view synthesis.
@@ 4.4 Results
Autoregressive Reconstruction We demonstrate the autoregressive reconstruction capability of our model in the second row of LABEL:fig:teaser. With only 4 input views, the model already produces reasonable 3D Gaussian reconstructions; as additional views arrive (8 and 32 views), both the rendering quality and scene coverage progressively improve. Additional examples of autoregressive reconstruction are provided on our project page. Table 2 also shows the quantitatively results of our autoregressive model, which constantly outperforms Long-LRM and remains competitive with, or superior to, optimization-based baselines.
@@ 4.5 Ablation Study
Optimizer and Losses We use Muon optimizer for its stabilty and robustness. Table 5 shows that use Muon as opitmizer can bring better results even on low resolution setting. It will bring better results on longer sequence (e.g. high resolution and more input views). Also, using depth and opacity as regularization can help reduce opaque Gaussians.
@@ B.1 Scene-level Training
We adopt a curriculum training strategy that progresses from low to high resolution, motivated by two main reasons. First, fast low-res pretraining enables the model to train with a large batch size with faster iteration time. Second, we found that even with pretrained TTT-LVSM [70] checkpoints at high-resolution (i.e. $960\times 540$), the model cannot predict reasonable Gaussians at the beginning iterations, leading to excessive GPU memory usage due to the rendering of a large number of Gaussians.
@@ Appendix C More results and Comparison
In Table 7, we show results where we combine our method with a few additional optimization steps. It demonstrates that the reconstructed model can be further improved with minimal optimization cost, surpassing both purely optimization-based methods and the previous state-of-the-art feed-forward method, Long-LRM, under the same post-optimization setup.